# Effective Potential and Topological Photon Spheres: A Novel Approach to Black Hole Parameter Classification

### v1

@@ 4.2 Photon Sphere and charged AdS black hole with perfect fluid dark matter(CPFDM)
As it is evident from table (2), the forbidden regions are located next to the allowed regions and for this black hole there is practically no region that is equivalent to the naked singularity in terms of total topological charges. As seen in Figs (7a) and (7b), the structure for two different values of $\lambda$ shows the behavior of a normal black hole with a total charge of -1 and an unstable photon sphere, which is confirmed with a maximum energy. In comparison with the results from the previous state, it appears that the addition of an electric field and the AdS radius to the action has diminished the impact of the PDFM term. This is because changes in the $\lambda$ parameter, which previously led to the emergence of energy minima beyond the event horizon and drove the model towards a naked singularity, now have no effect.
We became curious to know whether this effect was more a result of the presence of the electric field or the de Sitter radius. Consequently, we investigated a scenario that included only the electric field [49]. Our studies indicated that the outcomes of the new model are entirely consistent with the results of this scenario. It seems that the addition of an electric field to the model, within the normal range of electric charge, i.e., ( q/m ¡ 1 ) and not in extreme cases, eliminates the possibility of enhancing gravity by generating potential minima beyond the event horizon. However, since the graphical results with these obtained states were almost identical, we refrained from including them in the article to avoid confusion and excessive volume.
@@
Figure 7: The normal vector field $n$ in the $(r-\theta)$ plane. The photon spheres are located at $(r,\theta)=(2.352075997,1.57)$ for fig (7a) with respect to $(q=0.1,m=1,\lambda=0.2,l=1)$ and $(r,\theta)=(2.674176616,1.57)$, for fig (7b) with respect to $(q=0.1,m=1,\lambda=1.69,l=1)$, (7c): the topological potential H(r) for CPFDM black hole model
@@ 4.3 Photon Sphere and Euler-Heisenberg black hole
Figure 10: The normal vector field $n$ in the $(r-\theta)$ plane. The photon spheres are located at $(r,\theta)=(0.7031751797,1.57)$, $(r,\theta)=(2.483792443,1.57)$ for fig (10a) with respect to $(q=0.8,m=1,a=-1,l=1)$, fig (10b) the topological potential H(r) for Euler-Heisenberg black hole model
@@ 4.5 Photon Sphere and Black Holes With Nonlinear Electrodynamics(NLE)
The results of ”Gao’s” work on the black hole showed that to have three horizons, we need $\alpha_{2}<0$ It should be stated that our results confirm the above-mentioned limit for black hole behavior. Our results show that for $\alpha_{2}\leq 4.8148$ , the model shows the behavior of a regular black hole and has an unstable photon sphere ,Figs (6a). For the range between $18.10644.8148<\alpha_{2}<$ 18.1064 the behavior of the model is the naked singularity,Figs (6b) and for a limit greater than that, the system lacks any spherical photon structure, either stable or unstable, and practically, it will most likely lack a black hole structure table(6). Since the structure of the energy diagrams was similar to the past behaviors, we refrained from displaying them here.
@@
It seems that ”Gao” in this article [58] did not mention the allowed distances in the case of 4 horizons and selected the parameters based on the need and in an optional way. But we show that in this case, there can be meaningful and meaningless intervals for our choice table(7).
Unlike all previous shapes, in this case another parameter has been added to our equations. This means that, since we have only one condition to establish, so our restrictions are increased, and we have to give a value to one of the parameters and add it to our constants table, but this will not reduce the efficiency of our method.
Because, foremost, this condition protects us from being unaware of the parameters of the system and their limitations, and at least it is always possible to make a decision about the allowed ranges of one of the parameters with good accuracy.
Secondly, this optionally of parameter selection will leave our hands free to include the less effective parameter in our constants table according to the selection conditions. According to [58] for the metric function in the case of four-horizons, we will have
@@ 5 Conclusions
In this paper, we aim to provide a comprehensive analysis of ultra-compact objects with various gravitational structures by examining the existence and behavioral patterns of topological photon and anti-photon spheres. We approach this study by TTC and analyzing the extremums of the defined effective potential graph. The field diagrams, zero points, and effective potential graph are meticulously plotted, and the characteristics under study detailed in Tables 1-7.
Indeed, it should be noted that The ability to study the behavior of photon spheres solely through the metric function and a potential directly derived from it is a distinct advantage. Although, there are limitations due to spherical symmetry and the method’s applicability to structures beyond four dimensions. But, what distinguishes this work is its unique perspective on the necessity of a photon sphere for ultra-compact gravitational structures like black holes—a necessity confirmed by numerous empirical and theoretical studies [15, 16, 17, 18, 19, 20, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71]. Utilizing this necessity, which is essentially a bidirectional relationship between the photon sphere and the gravitational structure, can lead to several interesting implications:
@@
• As a preliminary outcome, the calculation of topological photon spheres can complement the metric function results and even serve as a suitable replacement for that in examining the behavior of gravitational structures. An unstable photon sphere only dominates the space of the system when an event horizon exists, therefore, the range in which the behavior of the model is a black hole, practically meets the basic requirements for WCCC. Additionally, the broader coverage of the topological photon sphere over the surrounding space of the gravitational structure makes it more distinguished than the metric function.
@@
• The most significant concept shaping the primary goal of the article is that we can easily classify the parameter space of the models under study based on the existence and location of topological photon and anti-photon spheres. This means we can precisely indicate which parameter range of the model corresponds to a black hole structure, a naked singularity, or even a model that mathematically lacks a physical and scientific response for its surrounding space, which we have termed the forbidden zone.[72]
@@
• Based on this study, a question has arisen for us that requires further detailed examination: If we slice the space-time of an ultra-compact body (a naked singularity) with a horizontal plane, it will have at least one ring. A stable photon sphere with a total charge of 1 or two distinct rings (one stable and the other unstable) with a total charge of 0. Considering that any photon trapped in a stable ring will never leave its orbit outwardly, So practically from this border towards the center of the structure, the area becomes invisible to us. Given the vital importance of the photon sphere in determining the shadow, this ring essentially acts as an observable boundary for us. Now, since according to the WCC conjecture, anything unpredictable must be hidden behind the event horizon, our question is whether hypothetically, the radius of the stable photon sphere can be considered a new hypothetical horizon for the space-time under study? What implications would this choice have for the structure of the ultra-compact object?
@@
• The other achievement of this method is that we can adjust the parameters of our system in such a way that the model under study is compelled to exhibit the desired behavior, or conversely, we can estimate the approximate range of model parameters based on the presence of the photon sphere in a specific region of space. For instance, in certain gravitational studies, we require a gravitational structure that adheres to specific principles and behaviors. In such cases, where we need the model’s parameters to fall within a specific range while simultaneously maintaining the structure in a defined form, such as a black hole, this method can play a significant role.
To elucidate further, in the model studying the Swampland Conjectures, there are hypotheses such as the Weak Gravity Conjecture (WGC) which necessitate that The structure under study, while being overly charged, still retains in the form of a black hole. In this work, we propose that for such scenarios, the topological behavior of the photon sphere can be utilized, and after setting the parameters as desired, the topological photon sphere can be examined to determine whether the optimal conditions can be achieved or not.
@@
Finally, it is imperative to highlight that during the formulation of a theory, parameters are added as necessitated by the theoretical framework. These parameters may conceptually possess a defined range from the outset or be devoid of any limitations. In this article, we have endeavored to address the question of what impact these theories, when incorporated into black hole action, will have on the behavior of the photon sphere for their permissible values. Furthermore, from the perspective of the photon sphere, whether these parameters, for all their previously allowed values, will result in the natural behavior of the model under study, or if new constraints must be imposed on the parameters in the context of black hole studies.

### v2

@@ 4.2 Photon Sphere and charged AdS black hole with perfect fluid dark matter(CPFDM)
As it is evident from table (2), the forbidden regions are located next to the allowed regions and for this black hole there is practically no region that is equivalent to the naked singularity in terms of total topological charges. As seen in Fig (7a), the structure shows the behavior of a normal black hole with a total charge of -1 and an unstable photon sphere, which is confirmed with a maximum energy Fig (7b). In comparison with the results from the previous state, it appears that the addition of an electric field and the AdS radius to the action has diminished the impact of the PDFM term. This is because changes in the $\lambda$ parameter, which previously led to the emergence of energy minima beyond the event horizon and drove the model towards a naked singularity, now have no effect.
@@
Figure 7: The normal vector field $n$ in the $(r-\theta)$ plane. The photon spheres are located at $(r,\theta)=(2.352075997,1.57)$ for Fig (7a) with respect to $(q=0.1,m=1,\lambda=0.2,l=1)$, (7b): the topological potential H(r) for CPFDM black hole model
@@ 4.3 Photon Sphere and Euler-Heisenberg black hole
Figure 10: The normal vector field $n$ in the $(r-\theta)$ plane. The photon spheres are located at $(r,\theta)=(0.7031751797,1.57)$, $(r,\theta)=(2.483792443,1.57)$ for Fig (10a) with respect to $(q=0.8,m=1,a=-1,l=1)$, Figs (10b),(10c) the topological potential H(r) for Euler-Heisenberg black hole model
@@ 4.5 Photon Sphere and Black Holes With Nonlinear Electrodynamics(NLE)
The results of ”Gao’s” work on the black hole showed that to have three horizons, we need $\alpha_{2}<0$ It should be stated that our results confirm the above-mentioned limit for black hole behavior. Our results show that for $\alpha_{2}\leq 4.8148$ , the model shows the behavior of a regular black hole and has an unstable photon sphere ,Figs (6a). For the range between $18.10644.8148<\alpha_{2}<$ 18.1064 the behavior of the model is the naked singularity,Figs (6b) and for a limit greater than that, the system lacks any spherical photon structure, either stable or unstable, and practically, it will most likely lack a black hole structure, table(6). Since the structure of the energy diagrams was similar to the past behaviors, we refrained from displaying them here.
@@
Unlike all previous shapes, in this case another parameter has been added to our equations. This means that, since we have only one condition to establish, so our restrictions are increased, and we have to give a value to one of the parameters and add it to our constants table.
According to [61] for the metric function in the case of four-horizons, we will have
@@ 5 Conclusions
Given the necessity of photon rings and photon spheres for gravitational structures in numerous studies[15, 16, 17, 18, 19, 20, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74], this paper, for the first time, adopted a novel approach by utilizing the existence and behavioral patterns of topological charges associated with photon and anti-photon spheres. We aimed to determine the parameter ranges and their impact on the behavior of ultra-compact objects with various gravitational structures.
To achieve this, we examined models with nonlinear fields or multiple horizons, often influenced by PDFM. We meticulously plotted field diagrams, zero points, total topological charges, and effective potential diagrams. Based on these results, we provided a detailed classification of the behavior and characteristics of each model in Tables 1-7.
Overall, it is essential to state that the bidirectional relationship between the photon sphere and the gravitational structure is an ideal advantage, and studying it from both perspectives can lead to several interesting outcomes:
@@
• As an initial result, calculating topological photon spheres can complement metric function results and even serve as a suitable alternative for examining the behavior of gravitational structures. An unstable photon sphere dominates the system’s space only when an event horizon exists. Therefore, the range in which the model behaves as a black hole essentially meets the basic requirements of the WCCC. Additionally, the broader spatial coverage of the topological photon sphere study method around the gravitational structure is a key point that distinguishes it from the metric function.
@@
• One of the main objectives of this paper is to easily classify the parameter space of the studied models based on the existence and topological location of photon and anti-photon spheres. This means we can precisely identify which parameter range of the model corresponds to a black hole structure or a naked singularity.[26]
@@
• According to this study, if we slice the spacetime of an ultra-compact object (a naked singularity) at the location of the photon sphere with a horizontal plane, there will be at least one ring. An unstable photon sphere with a total charge of -1 or several distinct rings (stable and unstable) with a total charge of 0 or +1. Given that any photon trapped in a stable ring never leaves its orbit, practically, from this boundary towards the center of the structure, it is an area considered invisible to us. Due to the critical importance of the photon sphere in determining the shadow, this ring essentially acts as a visible boundary for us. Now, since according to the WCCC conjecture, any unpredictable region must be hidden behind the event horizon, our question is whether, hypothetically, the radius of the stable photon sphere can be considered a new hypothetical horizon for the studied spacetime? What implications does this choice have for the structure of the ultra-compact object?
@@
• Another achievement of this method is that we can adjust the system parameters to force the studied model to exhibit the desired behavior or, conversely, estimate the approximate parameter ranges of the model based on the presence of the photon sphere in a specific region of space. For example, in some gravitational studies, we need a gravitational structure that adheres to specific principles and behaviors. In such cases, where we need the model parameters to fall within a specific range while simultaneously maintaining the structure as a defined black hole, this method can play a significant role. To further clarify, in the Swampland conjecture model study, hypotheses such as the Weak Gravity Conjecture (WGC) require that the studied structure, while being overly charged, still maintains its black hole structure. In this work, we propose that for such scenarios, the topological behavior of the photon sphere can be used, and after adjusting the parameters as desired, the topological photon sphere can be examined to determine whether optimal conditions can be achieved.
@@
• Finally, it is crucial to emphasize that during the formulation of a theory, parameters deemed necessary by the theoretical framework are added. These parameters may conceptually have a defined range from the outset or lack any limitations. In this paper, we have attempted to address the question of what impact these theories, when incorporated into the black hole model action, will have on the behavior of the photon sphere for their permissible values. Additionally, from the perspective of the photon sphere, whether these parameters, with all previous permissible values, will lead to the natural behavior of the studied model or if new constraints need to be applied to these parameters.